OpenStudy (anonymous):

What is the slope of a line that passes through the points (1, -2) and (-3, -6)? -2 -1 1 2

Parth (parthkohli):

There's a formula that the teachers want, or there's an easier method which you can do in MCQ.

OpenStudy (cwrw238):

slope = y2 - y1 / x2 - x1

OpenStudy (anonymous):

well im doing computer work so u dont have to sho work but there is a formula

OpenStudy (anonymous):

Slope is given by \[m = {y_2 - y_1 \over x_2 - x_1} = {-6 + 2 \over -3 -1} = {-4 \over -4} = 1\]

Parth (parthkohli):

You can try the easier method. x changes by -4. y changes by -4.

Parth (parthkohli):

Now just divide -4 and -4... and you get 1.

Parth (parthkohli):

Just sayin, but you could do that in your head.

OpenStudy (anonymous):

Which is what the formula means to say anyway, in the language of mathematics :)